3. a) Three messages A, B, C have probabilities 0.4, 0.3 and 0.3 respectively, of their

occurrences. If total 100 massages are sent out, calculate the total information (Itotal)

carried by them. Also calculate the entropy of the signal.


1
Expert's answer
2021-03-31T15:34:48-0400

Probability of message A, "p_1=0.4"

Probability of message A, "p_2=0.3"

Probability of message A, "p_3=0.3"


Total messsages L=100


Total information

"I_{total}=p_1Llog_2\dfrac{1}{p_1}+p_2Llog_2\dfrac{1}{p_2}+p_3Llog_2\dfrac{1}{p_3}"


"=(0.4)(100)log_2\dfrac{1}{0.4}+(0.3)(100)log_2\dfrac{1}{0.3}+(0.3)(100)log_2\dfrac{1}{0.3}"


"=40\times1.32+30\times 1.735+30\times 1.735"


"=52.8+52.06+52.06=156.92" bits


Entropy "=\dfrac{I_{total}}{L}=\dfrac{156.92}{100}=1.5692 bits"